# Evaluating Self-Supervised Learning in Medical Imaging: A Benchmark for Robustness, Generalizability, and Multi-Domain Impact

### v1

@@ 4.3 Generalizability
Table 4: Average Accuracy Drops in Transfer Performance: The accuracy drop percentages, relative to the mean in-domain performance using ResNet-50 with random initialization are averaged over all possible target datasets. The lowest accuracy drops are highlighted in green, indicating better generalizability. Due to space constraints, only five of the eight evaluated methods are displayed. However, the shown results correspond to the best-performing method out of the full set of eight candidates.
@@
We further examine the impact of supervised ImageNet1K initialization on model generalizability. Figure 8 illustrates the transfer performance difference for models pre-trained on DermaMNIST with random versus ImageNet1K initialization. In general, the results favor ImageNet1K initialization across the majority of evaluation scenarios. This trend holds across other datasets and self-supervised methods as well. To evaluate the overall impact of ImageNet1K initialization compared to random initialization, we conducted paired $t$-tests for each test dataset across all cross-dataset training combinations including the in-domain setting where test and train splits come from the same dataset. The results demonstrate statistically significant improvements $(p<0.05)$ across all datasets, with PneumoniaMNIST showing the most modest relative gain ($1.88\pm 3.28\%$) and OCTMNIST exhibiting the largest improvement ($13.22\pm 25.59\%$). Detailed quantitative results are provided in Appendix D.3. While previous studies have raised concerns about the efficacy of transfer learning from natural to medical images due to significant distributional differences [61], our findings show that continual pre-training [32, 67] improves both in-domain performance and generalization across diverse medical datasets.
Figure 8: Comparison of transfer performance for SSL methods pre-trained on DermaMNIST with ImageNet1K and random initializations (hatched). Only SimCLR, MoCo v3, and BYOL are chosen for better visualization.
While ResNet-50 achieved higher in-domain performance than ViT-Small in most datasets (8/11 when trained from scratch and 9/11 when starting from ImageNet1K weights), ViT-Small demonstrated superior cross-dataset generalization. Specifically, when averaged across different methods, ViT-Small’s performance consistently exceeded that of ResNet-50 by a significant margin in all datasets when trained from scratch. In the pretrained scenario, it outperformed ResNet-50 in 9 out of 11 datasets, with only marginal differences in the remaining two, as shown in Figure 9.
@@
Finally, we evaluate the generalizability of representations learned through multi-domain training. Figure 10 compares the cross-dataset performance of models trained on multi-domain datasets with models trained on their individual constituents. Training with similar domain datasets (Organ{A,C,S}) enhances cross-dataset transfer performance, suggesting that multi-domain training within a single modality may act as a regularizer. However, this benefit does not hold for mixed-modality combinations (Organ{A,S}PnePath), showing that multi-domain training benefits are modality-dependent (also see Appendix D.3).
@@ 5 Discussion & Conclusion
Despite concerns that supervised pre-training might yield less general representations due to label dependency [49, 62], ImageNet1K initialization enhances both in-domain and cross-dataset performance. However, this advantage does not consistently extend to OOD detection, as ImageNet1K initialization degrades performance for certain backbone and model combinations. Overall, supervised and self-supervised ImageNet1K initialization offer comparable performance, with no clear advantage for either.
@@
Choosing between these architectures involves key trade-offs: While ResNet-50 excels in in-domain classification, ViT-Small demonstrates superior performance in OOD detection and better generalization across datasets, achieving higher mean accuracies in all cross-dataset evaluations. Though these representations are more generalizable, ViT shows greater sensitivity to label scarcity, with performance dropping more steeply when training data is limited. Consistent with Matsoukas et al. [55], ImageNet1K initialization reduces performance gap, though the trade-offs remain.
@@
6 Acknowledgements
B.K. gratefully acknowledges partial support from the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) Master’s Scholarships.
@@ A Dataset Details
The MedMNIST dataset was initially introduced in a resolution of $28\times 28$ pixels and has since been expanded by MedMNIST+ to additionally include $64\times 64$, $128\times 128$, and $224\times 224$ resolutions. Due to limited GPU memory, we opted to use the $64\times 64$ resolution version of the dataset throughout our experiments. As lower resolutions are also shown to yield reasonable accuracies in previous research [38, 74], we do not sacrifice much by not using a larger resolution. Detailed information regarding each dataset, such as data source, domain, classification task type (including the number of classes), and publicly available data splits that correspond to our benchmark are provided in Table 5.
@@ C.1 Pre-Training
For training, we employ mini-batch gradient descent using the AdamW optimizer [53] and a cosine learning rate schedule [65] beginning with a $10$-epoch linear warm-up from $3\times 10^{-5}$. To ensure fairness, we conduct a grid search using learning rates of the form $3\times 10^{x}$, where $x\in\{-1,-2,-3,-4\}$, and weight decays of the form $1\times 10^{x}$, where $x\in\{-3,-4,-5\}$. We select the model with the best performance on the downstream validation set. The batch size is set to $256$, with coefficients for computing running averages of the gradient and its square as $\beta_{1}=0.9$ and $\beta_{2}=0.95$ respectively. We train the models for a total of 400 epochs.
For our experimental evaluation with self-supervised ImageNet1K pre-training, we employ four established methods: MoCo v3, SimCLR, BYOL, and DINO. The MoCo v3 weights were obtained from the mmselfsup library [18], specifically utilizing the model variant trained for $100$ epochs with a batch size of $4,096$ on the ImageNet1K dataset. For SimCLR, BYOL, and DINO, we leverage pre-trained weights from the LightlySSL framework [66]. These models were trained for 100 epochs, with varying batch sizes: 256 for SimCLR and BYOL, and 128 for DINO.
@@ C.2 Data Augmentations
All of the used SSL methods employ a sequence of image transformations to obtain different views from a sample. In our experiments, we have used augmentation configurations provided in the solo-learn library, as described below. Specifically, we have used the respective configurations of VICReg and ReSSL and the default asymmetric configuration for the other methods. Asymmetric here refers to the two contrasting views obtained from a single image sample. This default configuration in question applies the following augmentations (denoted as $T$) twice to obtain the two views:
Random Resized Crop ($T_{RRC}$): With probability $1$, a random crop of the image is extracted. The scale of the crop is uniformly sampled from the range $[0.2,1]$, while maintaining the aspect ratio. The cropped region is then resized to the original image dimensions.
Color Jitter ($T_{CJ}$): With probability $0.8$, the image undergoes random adjustments in brightness, contrast, saturation, and hue. The maximum intensity of these adjustments is set to $0.4$ for brightness and contrast, $0.2$ for saturation, and $0.1$ for hue.
Grayscale Conversion ($T_{Gray}$): With probability $0.2$, the RGB image is converted to grayscale by computing the luminance $L$ of each pixel using the formula $L=0.299R+0.587G+0.114B$ where $R$, $G$, and $B$ are the red, green, and blue channel values, respectively.
Gaussian Blur ($T_{GB}$): Applied with probability $1$, the image is blurred using a Gaussian filter. The standard deviation $\sigma$ of the Gaussian kernel is uniformly sampled from the range $[0.1,2]$.
Solarization ($T_{S}$): Solarization is applied asymmetrically across augmented views. For one view, it is applied with probability $0.2$, and for the other view, it is not applied (probability 00). The transformation inverts the pixel intensities for pixels with values $L\geq 128$ (assuming pixel intensities in the range $[0,255]$) using $L^{\prime}=255-L$ where $L$ is the original pixel intensity and $L^{\prime}$ is the transformed intensity.
Horizontal Flip ($T_{F}$): With probability $0.5$, the image is flipped horizontally.
The transformation sequence for ReSSL differs since it classifies the augmentations as weak and contrastive [80]. In the contrastive augmentations, the scale for $T_{RRC}$ is chosen uniformly from range $[0.08,1]$, $T_{GB}$ is applied with probability $0.5$ and $T_{S}$ is not applied. Rest of the augmentations are the same as above. For the weak augmentation the scale for $T_{RRC}$ is again chosen uniformly from range $[0.08,1]$, $T_{F}$ is applied with probability $0.5$ and there are no other augmentations applied.
VICReg augmentations also differ since they are applied symmetrically, that is, with the same parameters and probabilities for two different views. $T_{GB}$ is applied with probability $0.5$ and $T_{S}$ applied with probability $0.1$ for both views. The rest of the augmentations is the same as the default configuration.
@@ C.3 Linear Evaluation
We utilize the train, validation, and test set splits as provided by MedMNIST. All images are converted to RGB format and normalized. We use stochastic gradient descent with a step learning rate scheduler, decaying at epochs $[60,80]$. For a fair comparison, we conduct a grid search over learning rates {$0.1,0.01,0.001$} and weight decay values {$0.1,0$ 0.1 , 0.01}, selecting the best hyperparameters based on validation set performance. The test set is used only once at the end to evaluate the classifier with the optimal hyperparameters. We train the linear classifier for $100$ epochs. We report the mean and standard deviation of performance metrics on downstream tasks over five runs. Assuming a t-distribution, we calculate the confidence intervals as $\bar{x}\pm t\times\text{SE}$, where $\bar{x}$ represents the mean metric and SE is the standard error, calculated as $\text{SE}=\frac{\text{SD}}{\sqrt{n}}$ where $n$ is the number of trials and SD is the standard deviation.
@@ D.1.1 Linear Evaluation with All Labels
The in-domain performance of the self-supervised learning methods using ResNet-50 with random initialization is presented in Section 3 in Table 2. Additional results for ResNet-50 with ImageNet1K initialization, ViT-Small with random initialization, and ViT-Small with ImageNet1K initialization are provided in Tables 55, and 5 respectively. Similar to Table 2, Area Under the Curve (AUC) and accuracy (ACC) metrics are reported, and supervised learning results are provided for reference. The highest accuracy scores among the SSL methods are highlighted in green, and conversely, the lowest ones are highlighted in red.
These tables illustrate how initialization and architecture choices impact the performance of different SSL methods. Notably, DINO performs significantly better with ImageNet1K initialization on both ResNet-50 and ViT-Small, whereas it ranks among the worst with random initialization. Moreover, compared to the random initialization results for ResNet-50 in Table 2, SimCLR achieves notably higher accuracy with ViT-Small. Further analyses on the impact of initialization and backbone architecture are provided in Tables  6 and 7, respectively. Additionally, Figure 11 shows the performance differences between random and ImageNet1K initialization for each dataset and SSL method.
@@
Random vs ImageNet1K Initialization Performance Comparison: We analyze the impact of initialization strategies on self-supervised learning methods across various medical datasets in Table  6. The best scores in random initialization are underlined and the best scores in ImageNet1K initialization are bolded. MoCo v3 emerges as the best method under random initialization, achieving superior performance in 5 out of 11 datasets. The transition to ImageNet1K initialization shifts the performance landscape; DINO and BYOL show the most substantial improvements. DINO becomes the new leader, achieving the best accuracy in 3 out of 11 datasets. The results show that while ImageNet1K initialization generally enhances performance, the choice of self-supervised method should be tailored to specific medical imaging tasks, considering both the initialization strategy and the target domain’s characteristics.
@@
ResNet-50 vs. ViT-S Performance Comparison: To systematically assess the impact of architectural differences, we compare ResNet-50 and ViT-Small under both random and ImageNet1K initialization strategies. Table 7 presents the performance differential, computed as the accuracy of ResNet-50 minus that of ViT-Small. For visual clarity, we employ a color-coding scheme where red indicates superior ResNet-50 performance and blue denotes better ViT-Small performance. Cells are left uncolored when the absolute performance difference is negligible ($<0.01$ percentage points). ImageNet1K initialization generally reduces the performance gap between architectures, suggesting that pre-training helps mitigate architectural biases. Some datasets (e.g., OCT, OrganA) exhibit consistent architectural preferences regardless of initialization, while others (e.g., Blood, Breast) show initialization-dependent trends, where the preferred architecture shifts based on the initialization strategy.

### v2

@@ 4.3 Generalizability
Table 2: Average Accuracy Drops in Transfer Performance: The accuracy drop percentages, relative to the mean in-domain performance using ResNet-50 with random initialization are averaged over all possible target datasets. The lowest accuracy drops are highlighted in green, indicating better generalizability. Due to space constraints, we show only five of the eight evaluated methods, representing the best-performing candidates from the full set.
@@
[…]
ResNet-50 outperformed ViT-Small in most in-domain settings (8/11 datasets when trained from scratch and 9/11 when initialized with ImageNet weights). However, as Figure 8 demonstrates, the performance gap between these architectures narrows considerably when assessing generalization rather than in-domain performance. Notably, in the ImageNet1K pre-trained generalization scenario, ViT-S actually surpasses ResNet-50 on 8/11 datasets, reversing the previous trend. This suggests that ViT-S benefits more from transfer learning than ResNet-50 when applied to new datasets - supporting a similar finding [57].
@@
Finally, we evaluate the generalizability of representations learned through multi-domain training. Figure 9 compares the cross-dataset performance of models trained on multi-domain datasets with models trained on their individual constituents. Training with similar domain datasets (Organ{A,C,S}) enhances cross-dataset transfer performance, suggesting that multi-domain training within a single modality may act as a regularizer. However, this benefit does not hold for mixed-modality combinations (Organ{A,S}PnePath), showing that multi-domain training benefits are modality-dependent (see Appendix D.3).
@@ 5 Discussion & Conclusion
Self-supervised pretraining on medical images, initialized with ImageNet1K weights, significantly outperforms random initialization for in-domain downstream tasks, mirroring the established benefits previously observed in supervised learning setting [70]. Despite concerns that supervised pre-training might yield less general representations due to label dependency [51, 64], ImageNet1K initialization enhances both in-domain and cross-dataset performance. However, this advantage does not consistently extend to OOD detection, as ImageNet1K initialization degrades performance for certain backbone and model combinations. Overall, supervised and self-supervised ImageNet1K initialization offer comparable performance for in-domain accuracy and OOD detection with no clear advantage for either.
@@
Choosing between these architectures involves key trade-offs: While ResNet-50 excels in in-domain classification (outperforming ViT-Small in 8/11 datasets when trained from scratch and 9/11 with ImageNet initialization), ViT-Small demonstrates superior performance in OOD detection. Interestingly, when trained with ImageNet initialization and evaluated on generalization tasks, ViT-Small reverses the in-domain trend, surpassing ResNet-50 in 8/11 datasets. However, ViT shows greater sensitivity to label scarcity, with performance dropping more steeply when training data is limited. Consistent with Matsoukas et al. [57], ImageNet initialization reduces the performance gap between architectures, though these fundamental trade-offs remain.
@@
Extended Evaluation of Key Findings:
We conducted a series of short experiments using bigger architectures, higher resolution images, and an additional dataset AIROGS [20].
Larger backbones: Using RN-101 ($\sim$45M parameters) and ViT-B ($\sim$86M parameters), we conducted generalizability, OOD, and in-domain experiments on DermaMNIST, plus in-domain tests on BloodMNIST with MoCo v3 and DINO, yielding consistent results. In cross-dataset evaluations, ImageNet1K init. outperforms random init. ($\Delta{=}11.0\%$) and MoCo v3 surpasses DINO ($\Delta{=}10.5\%$) across all settings. For OOD detection, ViT shows superior performance ($\Delta\text{AUC}{=}0.026$, 95% CI:±0.019$\pm$ 0.019± 0.019) (Fig. 6) and MoCo v3 outperforms DINO when using RN-50 with random init. ($\Delta\text{AUC}{=}0.037$, 95% CI:±0.032$\pm$ 0.032± 0.032). For in-domain results (Fig. 1), ImageNet1K init. improves in-domain accuracy across methods and architectures ($\Delta{=}7.64\%$) considering both datasets.
Additional dataset: We trained MoCo v3 and BYOL on AIROGS, a large real-world dataset, containing ~113K color fundus images from ~60K patients. Aligned with our findings, ViT dominates RN-50 with random initialization ($\Delta\text{AUC}{=}0.05$, 95% CI:±0.041$\pm$ 0.041± 0.041) in OOD detection and MoCo v3 outperforms BYOL in cross-dataset evaluations ($\Delta{=}2.72\%$) and across all initializations and architectures in OOD detection ($\Delta\text{AUC}{=}0.081$, 95% CI:±0.041$\pm$ 0.041± 0.041).
Higher resolution: We trained MoCo v3, BYOL, DINO, SimCLR using 224×224 OrganAMNIST. Results were consistent with 64×64: For OOD detection, ViT-S outperforms RN50 ($\Delta$AUC=0.097, 95% CI:±0.073$\pm$ 0.073± 0.073), and performance hierarchy (Fig. 5) among RN50 models with random init. is: MoCo v3 (0.93) $>$ DINO (0.91) $>$ BYOL (0.90) $>$ SimCLR (0.82). Also, ImageNet1K init. leads to higher cross-dataset performance across methods and architectures ($\Delta{=}0.84\%$ in average). RN50 achieves better in-domain performance than ViT-S (Fig. 2), where the performance gap shrinks when switching from random init. ($\Delta{=}2.44\%$) to ImageNet1K initialization ($\Delta{=}1.66\%$).
Acknowledgements
The author acknowledges support from the Zuse School ELIZA, which provided essential funding that enabled their contributions to this work.
@@ A Dataset Details
The MedMNIST dataset was initially introduced in a resolution of $\times 28$ 28 pixels and has since been expanded by MedMNIST+ to additionally include $\times 64$ 64, $\times 128$ 128, and $\times 224$ 224 resolutions. Due to limited GPU memory, we opted to use the $\times 64$ 64 resolution version of the dataset throughout our experiments. As lower resolutions are also shown to yield reasonable accuracies in previous research [39, 76], we do not sacrifice much by not using a larger resolution. Detailed information regarding each dataset, such as data source, domain, classification task type (including the number of classes), and publicly available data splits that correspond to our benchmark are provided in Table 3.
@@ C.1 Pre-Training
For training, we employ mini-batch gradient descent using the AdamW optimizer [55] and a cosine learning rate schedule [67] beginning with a 101010-epoch linear warm-up from $3\times 10^{-5}$. To ensure fairness, we conduct a grid search using learning rates of the form $3\times 10^{x}$, where $x\in\{-1,-2,-3,-4\}$, and weight decays of the form $1\times 10^{x}$, where $x\in\{-3,-4,-5\}$. We select the model with the best performance on the downstream validation set. The batch size is set to 256256256, with coefficients for computing running averages of the gradient and its square as $\beta_{1}=0.9$ and $\beta_{2}=0.95$ respectively. We train the models for a total of 400 epochs.
For our experimental evaluation with self-supervised ImageNet1K pre-training, we employ four established methods: MoCo v3, SimCLR, BYOL, and DINO. The MoCo v3 weights were obtained from the mmselfsup library [18], specifically utilizing the model variant trained for 100100100 epochs with a batch size of 4,0964,0964 , 096 on the ImageNet1K dataset. For SimCLR, BYOL, and DINO, we leverage pre-trained weights from the LightlySSL framework [68]. These models were trained for 100 epochs, with varying batch sizes: 256 for SimCLR and BYOL, and 128 for DINO.
@@ C.2 Data Augmentations
Azizi et al. [7] work on different medical modalities, including dermatology photography, fundus imaging, digital pathology, chest radiography, and mammography. Similar to their default contrastive pretraining setting, we utilized random cropping, random color distortion (brightness, contrast, saturation, and hue changes), and random Gaussian blur, along with random horizontal flipping as the data augmentation strategy.
@@ C.3 Linear Evaluation
We utilize the train, validation, and test set splits as provided by MedMNIST. All images are converted to RGB format and normalized. We use stochastic gradient descent with a step learning rate scheduler, decaying at epochs [60,80][60,80][ 60 , 80 ]. For a fair comparison, we conduct a grid search over learning rates {0.1,0.01,0.0010.1,0.01,0.0010.1 , 0.01 , 0.001} and weight decay values {$0.1,0$ 0.1 , 0.01}, selecting the best hyperparameters based on validation set performance. The test set is used only once at the end to evaluate the classifier with the optimal hyperparameters. We train the linear classifier for 100100100 epochs. We report the mean and standard deviation of performance metrics on downstream tasks over five runs. Assuming a t-distribution, we calculate the confidence intervals as $\bar{x}\pm t\times\text{SE}$, where $\bar{x}$ represents the mean metric and SE is the standard error, calculated as $\text{SE}=\frac{\text{SD}}{\sqrt{n}}$ where $n$ is the number of trials and SD is the standard deviation.
@@ D.1.1 Linear Evaluation with All Labels
The in-domain performance of the self-supervised learning methods using ResNet-50 with random initialization is presented in Section 3 in Table 4. Additional results for ResNet-50 with ImageNet1K initialization, ViT-Small with random initialization, and ViT-Small with ImageNet1K initialization are provided in Tables 44, and 4 respectively. Similar to Table 4, Area Under the Curve (AUC) and accuracy (ACC) metrics are reported, and supervised learning results are provided for reference. The highest accuracy scores among the SSL methods are highlighted in green, and conversely, the lowest ones are highlighted in red.
These tables illustrate how initialization and architecture choices impact the performance of different SSL methods. Notably, DINO performs significantly better with ImageNet1K initialization on both ResNet-50 and ViT-Small, whereas it ranks among the worst with random initialization. Moreover, compared to the random initialization results for ResNet-50 in Table 4, SimCLR achieves notably higher accuracy with ViT-Small. Further analyses on the impact of initialization and backbone architecture are provided in Tables  5 and 6, respectively. Additionally, Figure 10 shows the performance differences between random and ImageNet1K initialization for each dataset and SSL method.
Table 4: In-domain performance of SSL methods on a randomly initialized ResNet-50 backbone, reported as Area Under the Curve (AUC) and Accuracy (ACC). Supervised learning results with random and ImageNet1K initialization are included for comparison..
@@
Random vs. ImageNet1K Initialization Performance Comparison: We analyze the impact of initialization strategies on self-supervised learning methods across various medical datasets in Table  5. The best scores in random initialization are underlined and the best scores in ImageNet1K initialization are bolded. MoCo v3 emerges as the best method under random initialization, achieving superior performance in 5 out of 11 datasets. The transition to ImageNet1K initialization shifts the performance landscape; DINO and BYOL show the most substantial improvements. DINO becomes the new leader, achieving the best accuracy in 3 out of 11 datasets. The results show that while ImageNet1K initialization generally enhances performance, the choice of self-supervised method should be tailored to specific medical imaging tasks, considering both the initialization strategy and the target domain’s characteristics.
@@
ResNet-50 vs. ViT-S Performance Comparison: To systematically assess the impact of architectural differences, we compare ResNet-50 and ViT-Small under both random and ImageNet1K initialization strategies. Table 6 presents the performance differential, computed as the accuracy of ResNet-50 minus that of ViT-Small. For visual clarity, we employ a color-coding scheme where red indicates superior ResNet-50 performance and blue denotes better ViT-Small performance. Cells are left uncolored when the absolute performance difference is negligible (<0.01<0.01< 0.01 percentage points). ImageNet1K initialization generally reduces the performance gap between architectures, suggesting that pre-training helps mitigate architectural biases. Some datasets (e.g., OCT, OrganA) exhibit consistent architectural preferences regardless of initialization, while others (e.g., Blood, Breast) show initialization-dependent trends, where the preferred architecture shifts based on the initialization strategy.
@@
Self-Supervised ImageNet1K Weights vs. Supervised ImageNet1K Weights: Supervised ImageNet1K initialization outperformed self-supervised in 24/44 (54.5%) dataset-method combinations, while self-supervised performed better in 20/44 (45.5%). The difference between methods ranged from -9.25% (favoring self-supervised) to +12.35% (favoring supervised), with a median absolute difference of 1.52%. A paired t-test across all dataset-method combinations confirmed no statistically significant difference between supervised and self-supervised ImageNet1K initialization for in-domain performance (t(43) = 1.040, p = 0.3043, 95% CI: [-0.52%, 1.64%]), though both initialization techniques substantially outperformed random initialization.
Dataset MoCo v3 SimCLR DINO BYOL
Sup. SSL Sup. SSL Sup. SSL Sup. SSL
Path 92.84 93.52 92.81 92.52 94.05 93.85 93.85 93.43
Derma 78.50 77.29 77.29 76.66 78.63 80.21 78.99 77.12
OCT 82.94 81.48 80.98 75.46 80.64 80.10 82.44 80.56
Pneumonia 92.18 94.36 94.81 93.78 93.85 92.12 93.88 91.47
Retina 50.05 51.40 51.15 52.35 46.15 55.40 54.10 41.75
Breast 88.59 84.62 87.18 90.00 88.08 89.49 88.46 85.77
Blood 97.90 98.12 97.88 97.99 98.35 98.22 98.08 98.29
Tissue 59.09 63.27 60.65 62.90 59.43 47.21 59.58 61.23
OrganA 91.85 91.70 87.69 91.02 95.13 93.43 92.70 91.58
OrganC 90.47 91.26 91.47 89.58 91.51 91.81 89.38 91.01
OrganS 79.69 77.39 77.72 78.91 78.48 79.75 80.63 76.18